OpenStudy (anonymous):

Solve for x: 10x + 5 > 5 A. x = 10 B. x > 0 C. x < -1 D. x > 20

OpenStudy (anonymous):

can anyone help

OpenStudy (icalibear):

Yes.

OpenStudy (icalibear):

Basically what the question is asking is what is going to be bigger than 5 once you solve the equation, so can you eliminate any answers that are not bigger than 5?

OpenStudy (anonymous):

c?

OpenStudy (icalibear):

10(-1) + 5 = ?

OpenStudy (icalibear):

Can you answer that?

OpenStudy (anonymous):

14?

OpenStudy (icalibear):

So that's bigger than 5.

OpenStudy (icalibear):

You know now that that is a possible answer.

OpenStudy (hunus):

It's asking for what values of x is the left side bigger than the right side?

OpenStudy (icalibear):

Basically. Try plugging in numbers and see which one works.

OpenStudy (anonymous):

10x + 5 > 5 (subtract 5 from both sides) 10x + 5 - 5 > 5 - 5 10x = 0 x = 0/10 = 0

OpenStudy (icalibear):

Does that work for the equation?

OpenStudy (anonymous):

It doesn't seem to fit the equation...but that is how to do an inequality. You solve an inequality kind of like an equation. Except there are some rules... one of which is you can subtract/add anything to both sides as long as they are the same number...doesn't matter if it is negative or positive. (which I did) If this is done it does not change the inequality sign.Also if you divide by a positive number the inequality sign does not change. I can't see anything that I did wrong, however, the inequality doesn't seem to make sense.... 10x + 5 > 5 10 (0) + 5 > 5 0 + 5 > 5 5 > 5 (doesn't make sense ???)

OpenStudy (anonymous):

It ends up with 10x > 0 so x > 0!